# MOHPER: Multi-objective Hyperparameter Optimization Framework for E-commerce Retrieval System

### v2

@@ 2.2. Joint Optimization in E-Commerce Context
    Click-Through Conversion Rate (CTCVR): #$\#\text{purchases}\ /\#\text{imp.}$.
@@
    Click-Through Rate (CTR): #$\#\text{clicks}\ /\ \#\text{imp.}$.
@@
    Click-Through Cart-Adding Rate (CTCAR): #$\#\text{carts}\ /\ \#\text{imp.}$.
[…]
@@ 3.1.3. Hyperparamter Optimization Framework
Method CTR nDCG@20 CTR nDCG@100 CTR prec@100 CTR avg metrics CTCVR nDCG@20 CTCVR nDCG@100 CTCVR prec@100 CTCVR avg metrics Overall avg metrics
@@
Method CTR nDCG@20 CTR nDCG@100 CTR prec@100 CTR avg metrics CTCVR nDCG@20 CTCVR nDCG@100 CTCVR prec@100 CTCVR avg metrics Overall avg metrics
@@ 3.1.5. Samplers
TPE (Weighted-sum Based) $\mathcal{O}(kdn+kQL+n\log n)$ Sampling: $kdn$; Retrieval + Eval: $kQL$; Quantile: $n\log n$
TPE (Separate Multi-objective) $\mathcal{O}(kdnm+kQ(L+m)+nm\log n)$ Sampling: $kdnm$; Retrieval: $kQL$; Eval: $kQm$; Quantile: $nm\log n$
NSGA-III (Weighted-sum Based) $\mathcal{O}(gNd+gNQL)$ Sampling: $gNd$; Retrieval + Eval: $gNQL$
@@ 3.2. Experimental Results and Discussion
Exp. CTR gain(%) CTCAR gain(%) CTCVR gain(%) Click/Q gain(%) Cart/Q gain(%) Purchase/Q gain(%)
on SERP 0.95%*** 0.75%*** 0.30% 1.23%*** 0.98%* 0.75%*
on CPLP -0.51%*** 1.16%*** 1.10%* -0.22%** 1.45%** 1.62%*

### v3

@@ 2.2. Joint Optimization in E-Commerce Context
    Click-Through Conversion Rate (CTCVR): $\text{purchases}\ /\#\text{imp.}$
@@
    Click-Through Rate (CTR): $\text{clicks}\ /\ \#\text{imp.}$
@@
    Click-Through Cart-Adding Rate (CTCAR): $\text{carts}\ /\ \#\text{imp.}$
Here, $\text{imp.}$ represents the number of item impressions in retrievals. CTCVR is the primary objective, with CTR, and CTCAR serving as auxiliary objectives to mitigate selection bias and sparsity (Figure 3) (Ma et al., 2018; Karmaker Santu et al., 2017). Specifically, For any query $q$ and trial $t$, performance metrics $z_{t,q,(\cdot)}$ (e.g. precision@K, nDCG@K, mAP@K, and recall@K) are computed when ground truth labels $y_{q,(\cdot)}$ are available and the top-K items are retrieved ($\hat{y}_{h_{t},q,(\cdot)}$). Labels for these items are derived from objective-specific statistics (e.g. $imp_{q,i}$, $click_{q,i}$, $purchase_{q,i}$), with thresholding for positive/negative scoring at the corresponding objective. Also, while naive statistical rates are available, filtering based on impression counts and Bayesian smoothing for rare events can further refine these scores (Wang et al., 2011). Following metric calculation $z_{t,m,q}$, metrics are aggregated across queries, and MOHPO proceeds with multi-objective tuning to enhance engagement, conversion efficiency, and purchase intent, advancing retrieval performance in e-commerce.
@@ 3.1.3. Hyperparamter Optimization Framework
Method
CTR
nDCG@20
CTR
nDCG@100
CTR
prec@100
CTR
avg metrics
CTCVR
nDCG@20
CTCVR
nDCG@100
CTCVR
prec@100
CTCVR
avg metrics
Overall
avg metrics
@@
Method
CTR
nDCG@20
CTR
nDCG@100
CTR
prec@100
CTR
avg metrics
CTCVR
nDCG@20
CTCVR
nDCG@100
CTCVR
prec@100
CTCVR
avg metrics
Overall
avg metrics
@@ 3.1.5. Samplers
TPE (Weighted-sum Based) $\mathcal{O}(kdn+kQL+n\log n)$ Sampling: kdnkdn; Retrieval + Eval: $kQL$; Quantile: $n\log n$
TPE (Separate Multi-objective) $\mathcal{O}(kdnm+kQ(L+m)+nm\log n)$ Sampling: kdnmkdnm; Retrieval: $kQL$; Eval: $kQm$; Quantile: $nm\log n$
NSGA-III (Weighted-sum Based) $\mathcal{O}(gNd+gNQL)$ Sampling: $gNd$; Retrieval + Eval: gNQLgNQL
@@ 3.2. Experimental Results and Discussion
Exp.
CTR
gain(%)
CTCAR
gain(%)
CTCVR
gain(%)
Click/Q
gain(%)
Cart/Q
gain(%)
Purchase/Q
gain(%)
on SERP
0.95%*** 0.75%*** 0.30% 1.23%*** 0.98%* 0.75%*
on CPLP
-0.51%*** 1.16%*** 1.10%* -0.22%** 1.45%** 1.62%*